OpenStudy (anonymous):

what is the slope intercept equation of a line that has a slope of -2 and a y intercept of 3

OpenStudy (anonymous):

y = -2x+3

OpenStudy (anonymous):

so the slope intercept equation is y=mx+b?

OpenStudy (anonymous):

slope intercept form is y=mx+b where m is the slope, b is the y-intercept, and (x,y) are the coordinates of any point on the line. m=slope=-2 b=y-intercept=3 so y=-2x+3

OpenStudy (anonymous):

The slope intercept equation is y = mx + b m is the slope and b is the y intercept